# Beyond algorithm hyperparameters: on preprocessing hyperparameters and associated pitfalls in machine learning applications

### v1

@@ 5.3 Results
The optimistic bias of the apparent error is most pronounced in analysis settings where the preprocessing HPs $\boldsymbol{\lambda}_{P}$ are tuned manually (II-manual-P). This is not surprising, as this procedure specifically selects the HP values that optimize the apparent error. Here, the bias is largest when the smaller sample size and the CART algorithm are used for model generation, resulting in a median difference of 7.39 for RMSE and 0.253 for $R^{2}$. Note that while the absolute values of $\widehat{\mathrm{PE}}_{\text{train}}$ still do not indicate good predictive performance in these analysis settings (see Figure S2), the median $R^{2}$ values resulting from the CART algorithm (0.234 and 0.176 for the two sample sizes) are comparable to the prediction errors reported for the Australian and UK decision tree models (0.17 and 0.27), which were generally deemed viable (Eagar et al.,, 2004; Murtagh et al.,, 2023). Regarding the selected HPs, particularly for $\lambda_{ipos}$ (which specifies how the IPOS score is calculated) and $\lambda_{ca}$ (which determines how “cannot assess” values in IPOS features are handled), alternative values are frequently chosen instead of the defaults (see Figures S3a to S6a). This suggests that these alternative values may present a high potential for overfitting, thereby improving the apparent error.
@@
The lowest optimistic bias for the apparent error is observed in the analysis settings where either only $\boldsymbol{\lambda}_{A}$ (II-automated-A) or both $\boldsymbol{\lambda}_{P}$ and $\boldsymbol{\lambda}_{A}$ (II-automated-PA) are tuned automatically, with the largest median difference being 3.22 for RMSE and 0.035 for $R^{2}$. This is not surprising, as in these procedures, all HPs are selected based on their associated CV error estimate, rather than the apparent error. Notably, across all analysis settings, the HP values for $\boldsymbol{\lambda}_{P}$ selected by the II-automated-PA procedure differ from those chosen by the II-manual-P and II-combined-PA procedures (see Figures S3a to S6a).
@@
In the analysis settings without HP tuning, the $R^{2}$ differences exhibit a negative median difference, with the median difference closest to zero, -0.059, observed for the smaller sample size combined with the CART algorithm. This pessimistic bias is an expected result, as CV evaluates models trained on fewer observations than the final prediction model (see Section 3.2.1). In contrast to $R^{2}$, the prediction error differences for RMSE in the analysis settings without tuning are mostly positive. Although the median differences are small (with the largest median difference being 2.32 in the analysis setting where both the smaller sample size and the CART algorithm are considered), the overall distribution of the prediction error differences in each setting suggests the presence of an optimistic bias. This finding is unexpected, as prediction errors estimated by CV in a setting where no HPs are tuned should not exhibit an optimistic bias, but rather a pessimistic bias (as observed for $R^{2}$). However, this can be attributed to the fact that both $\widehat{\mathrm{PE}}_{\text{train}}$ based on CV and $\widehat{\mathrm{PE}}_{\text{new}}$ are affected by data leakage stemming from a violation of the assumption that all observations are independently drawn from the same distribution (see Section 2.4.2 and Supplementary Section A). This type of data leakage is distinct from the leakage caused by the overlap between the data used for model generation and evaluation, which is the primary focus of this paper. Specifically, the COMPANION data set exhibits a clustering structure that is not accounted for during the split into $\mathcal{D}_{\text{train}}$ and $\mathcal{D_{\text{new}}}$ or during the creation of CV splits on $\mathcal{D}_{\text{train}}$, resulting in a potential optimistic bias for both $\widehat{\mathrm{PE}}_{\text{new}}$ (due to the initial split) and $\widehat{\mathrm{PE}}_{\text{train}}$ (due to the CV splits). As $\widehat{\mathrm{PE}}_{\text{train}}$ is also subject to a larger clustering-induced optimistic bias than $\widehat{\mathrm{PE}}_{\text{new}}$, the bias does not cancel out when taking their difference and is therefore evident in Figure 5. Notably, the different levels of clustering-induced optimistic bias in $\widehat{\mathrm{PE}}_{\text{train}}$ and $\widehat{\mathrm{PE}}_{\text{new}}$ appear to have less impact on $R^{2}$, where, as described above, the prediction error differences are mostly negative. Further details on the impact of the clustering structure on the results, including an explanation of why it was not considered when performing the splits, are provided in Supplementary Section B.5.
@@
The additional source of optimistic bias introduced by the clustering structure of the data is also relevant when interpreting the prediction error differences in the analysis settings with HP tuning. While our primary focus here is on overlap-induced data leakage that arises since the observations used for the CV-based error estimation have already been seen during HP tuning (thus hindering the detection of potential overtuning), we have to consider that any observed optimistic bias may as well stem from clustering-induced data leakage. Consequently, we compare the prediction error differences in analysis settings with HP tuning to those in settings without tuning (where only clustering-induced data leakage is present), rather than directly comparing them to zero. Based on this assessment, the impact of overlap-induced data leakage on $\widehat{\mathrm{PE}}_{\text{train}}$ appears to be limited. This is particularly true for RMSE, where the CV error differences are generally comparable to those resulting from the I-no tuning procedure. For $R^{2}$, the median differences tend to be closer to zero compared to the I-no tuning procedure. In some analysis settings involving the smaller sample size and the CART algorithm, there is even a positive median difference (with the largest median difference of 0.018 observed in the setting where II-automated-PA is used in combination with the smaller sample size and the CART algorithm). Consequently, there appears to be a small overtuning effect that is not detected by the CV error due to overlap-induced data leakage. However, the median differences are too close to zero, and the variation within each analysis setting is too large to definitively determine which bias ultimately predominates, i.e. whether the CV error is overall optimistic or pessimistic in these settings.
@@
To summarize, the choice of model generation and evaluation procedure generally affects the difference between the prediction error estimates derived from available data and new data. As expected, when the evaluation procedure is based on the apparent error, the resulting estimate exhibits an optimistic bias, which varies depending on the model generation procedure. As likewise expected, the simple CV error is less optimistic than the apparent error, while the nested CV error is even less optimistic. The corresponding prediction error differences are less variable across model generation procedures compared to the apparent error. For simple CV, this indicates that, in the considered experimental setup, the tuning procedures do not introduce relevant overtuning effects on error estimation. Instead, the main source of bias for simple CV is either the clustering-induced optimistic bias (or more precisely, the different bias level relative to $\widehat{\mathrm{PE}}_{\text{new}}$) or the pessimistic bias arising from the use of fewer observations during evaluation. This also holds true for the nested CV error.
@@ 6 Discussion and conclusion
To provide different examples of model generation and evaluation procedures in the context of HP tuning, and to examine their impact on the difference between prediction error estimates from available and new data, we conducted an illustrative study using a real-world prediction problem from palliative care medicine. Although both the apparent error and CV error can, in theory, be optimistically biased when HPs are tuned, this was consistently true only for the apparent error (with the highest optimistic bias occurring in analysis settings that imitated manual tuning of preprocessing HPs without considering algorithm HPs). In contrast, the prediction error differences for the CV error appeared not to be considerably compromised by data leakage, as these differences were comparable to the analysis settings without HP tuning.
@@
Based on these conceptual and empirical insights, it is clear that to ensure HP tuning becomes a benefit rather than a pitfall, applied ML users must take care throughout the entire model development process. First, they should thoroughly consider which HPs (including preprocessing HPs) are to be tuned and which are not. An adequate tuning procedure should then be specified that fits the specific prediction problem. Unfortunately, this is typically non-trivial, as it depends on various factors such as sample size and the specific HPs to be tuned. More research is needed to better guide users in this respect (see Bischl et al.,, 2023 for an overview of current recommendations). In general, it is recommended to use automated tuning procedures instead of manual ones (see again Bischl et al.,, 2023 for automated tuning implementations in R and Python). If automated tuning is not feasible, users should at least ensure that the manual tuning procedure is error-free, reproducible, and resampling-based. For model evaluation, only two evaluation procedures are guaranteed to be unaffected by data leakage caused by HP tuning: (i) nested resampling (if the entire data set is used for model generation), or (ii) a permanent (outer) holdout (if only a subset of the available data is used for model generation). However, similar to the tuning procedure, there is a lack of guidance on how to choose between these approaches and how to specify them (e.g., which resampling methods to use for nested resampling). Although simple resampling may turn out to be a viable option in some applications (including our example), this can generally not be known in advance. Therefore, we discourage its use in settings involving HP tuning, as well as any other evaluation procedures that could result in data leakage.
Regardless of how model generation and evaluation are performed, it is essential that they and all other relevant details (e.g., the complete learning pipeline and its HPs) are transparently reported in both code and text form. For this purpose, users may rely on checklists such as REFORMS (Kapoor et al.,, 2024; intended for all applied research fields using ML) or TRIPOD-AI (Collins et al., 2024b, ; intended for clinical prediction models). While transparency does not imply correctness, it allows readers to identify potential issues such as data leakage, and to critically interpret the claimed model performance. Moreover, it emphasizes the existence and importance of preprocessing and its HPs, while the current lack of transparency can create the impression that the data were not preprocessed at all or that no alternative preprocessing options were explored. To further enhance transparency and encourage applied ML users to be more intentional about their choices, it is also possible to preregister the entire model development process, for example, by using the template proposed by Hofman et al., (2023).
@@
This work was supported by the German Research Foundation (BO3139/9-1, BO3139/7) to ALB. The authors of this work take full responsibilities for its content.
@@ Appendix A Other leakage types
As stated in Section 2.4.2, Kapoor and Narayanan, (2023) identify three general types of data leakage, which may arise from: (i) overlap between the data used for model generation and evaluation, (ii) violation of the assumption that all observations are independently drawn from the same distribution, or (iii) use of illegitimate features. While our paper primarily addresses overlap-induced data leakage, we will now provide additional details on the other two types.
@@
Even with a strict separation between the data used for model generation and evaluation, achieved through the use of resampling methods, data leakage can still occur if the assumption that all observations in $\mathcal{D}_{\text{train}}$ are independently drawn from the same distribution is violated. This assumption, also known as the i.i.d. assumption, was stated in Section 2.1. Non-i.i.d. settings may, for example, arise when $\mathcal{D}_{\text{train}}$ is a clustered data set, i.e. when the observations originate from different clusters (e.g., study centers). Observations within clusters are typically more similar than observations between clusters, where similarity can refer to both the feature vector $\boldsymbol{x}^{(i)}$ or the outcome $y^{(i)}$ (Hornung et al.,, 2023). If the prediction model is intended to be applied to observations from other clusters than those present in $\mathcal{D}_{\text{train}}$ in the future, resampling methods that are based on random sampling (i.e. ignoring the cluster structure) will be optimistically biased since in each resampling iteration, the observations in $\mathcal{D}_{\text{test}}$ are more similar to $\mathcal{D}^{\prime}_{\text{train}}$ than observations originating from new clusters (Hornung et al.,, 2023; Kapoor and Narayanan,, 2023; Rosenblatt et al.,, 2024). Although the level of optimistic bias depends on the specific clustering structure (e.g., cluster size and correlation within clusters), it is generally recommended to perform grouped resampling at cluster level, where all observations in a cluster are either assigned to $\mathcal{D}^{\prime}_{\text{train}}$ or $\mathcal{D}_{\text{test}}$ in each resampling iteration (Hornung et al.,, 2023; Bischl et al.,, 2023). In the context of healthcare research, this type of resampling is referred to as internal-external validation (Collins et al., 2024a, ; Debray et al.,, 2023). For other examples of non-i.i.d. settings and corresponding resampling methods, see Hornung et al., (2023) and the references therein.
@@
In Setting II, where resampling is typically used for both model generation (tuning) and evaluation, data leakage due to the violation of the i.i.d. assumption biases the prediction error estimate of the final model only when the non-i.i.d. data structure is ignored during model evaluation. This occurs specifically in the outer resampling loop of nested resampling (for $\mathcal{D}_{\text{train}}=\mathcal{D}$) or in the permanent outer holdout (for $\mathcal{D}_{\text{train}}\subset\mathcal{D}$). However, it is recommended to also take into account the non-i.i.d. data structure during tuning, both for the final prediction model and, if nested resampling is used, within the inner resampling loop, to ensure consistency (Hornung et al.,, 2023).
@@ A.2 Use of illegitimate features
If $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ include features that are generally not available for new observations to which the model will be applied in practice, these features can be considered illegitimate, and if included in the final prediction model, constitute another type of data leakage. An example raised by Kapoor and Narayanan, (2023) is the use of anti-hypertensive drugs as a feature for predicting hypertension. Note that this type of data leakage is conceptually different from the other two types, as it stems from a design issue that is independent of the model evaluation procedure.
@@ B.1 Descriptive statistics
Table S1: Distribution of the outcome variable and features in the COMPANION data set after applying the initial preprocessing steps (described in Supplementary Section B.2.1). In addition, two preprocessing steps from the learning pipeline $\mathcal{I}$ (see Sections 5.2.2 and B.2.2) have been performed: the correction of costs and the aggregation of the IPOS score (default version).
$n=$1,449

### v2

@@ 5.3 Results
The optimistic bias of the apparent error is most pronounced in analysis settings where the preprocessing HPs $\boldsymbol{\lambda}_{P}$ are tuned manually (II-manual-P). This is not surprising, as this procedure specifically selects the HP values that optimize the apparent error. Here, the bias is largest when the smaller sample size and the CART algorithm are used for model generation, resulting in a median difference of 7.39 for RMSE and 0.253 for $R^{2}$. Note that while the absolute values of $\widehat{\mathrm{PE}}_{\text{train}}$ still do not indicate good predictive performance in these analysis settings (see Figure S2), the median $R^{2}$ values resulting from the CART algorithm (0.234 and 0.176 for the two sample sizes) are comparable to the prediction errors reported for the Australian and UK decision tree models (0.17 and 0.27), which were generally deemed viable [92, 116]. Regarding the selected HPs, particularly for $\lambda_{ipos}$ (which specifies how the IPOS score is calculated) and $\lambda_{ca}$ (which determines how “cannot assess” values in IPOS features are handled), alternative values are frequently chosen instead of the defaults (see Figures S3a to S6a). This suggests that these alternative values may present a high potential for overfitting, thereby improving the apparent error.
@@
The lowest optimistic bias for the apparent error is observed in the analysis settings where either only $\boldsymbol{\lambda}_{A}$ (II-automated-A) or both $\boldsymbol{\lambda}_{P}$ and $\boldsymbol{\lambda}_{A}$ (II-automated-PA) are tuned automatically, with the largest median difference being 3.22 for RMSE and 0.035 for $R^{2}$. This is not surprising, as in these procedures, all HPs are selected based on their associated CV error estimate rather than the apparent error. Notably, across all analysis settings, the HP values for $\boldsymbol{\lambda}_{P}$ selected by the II-automated-PA procedure differ from those chosen by the II-manual-P and II-combined-PA procedures (see Figures S3a to S6a).
@@
In the analysis settings without HP tuning, the $R^{2}$ differences exhibit a negative median difference, with the median difference closest to zero, -0.059, observed for the smaller sample size combined with the CART algorithm. This pessimistic bias is an expected result, as CV evaluates models trained on fewer observations than the final prediction model (see Section 3.2.1). In contrast to $R^{2}$, the prediction error differences for RMSE in the analysis settings without tuning are mostly positive. Although the median differences are small (with the largest median difference being 2.32 in the analysis setting where both the smaller sample size and the CART algorithm are considered), the overall distribution of the prediction error differences in each setting suggests the presence of an optimistic bias. This finding is unexpected, as prediction errors estimated by CV in a setting where no HPs are tuned should not exhibit an optimistic bias but rather a pessimistic bias (as observed for $R^{2}$). However, this can be attributed to the fact that both $\widehat{\mathrm{PE}}_{\text{train}}$ based on CV and $\widehat{\mathrm{PE}}_{\text{new}}$ are affected by data leakage stemming from a violation of the assumption that all observations are independently drawn from the same distribution (see Section 2.4.2 and Supplementary Section A). This type of data leakage is distinct from the leakage caused by the overlap between the data used for model generation and evaluation, which is the primary focus of this paper. Specifically, the COMPANION data set exhibits a clustering structure that is not accounted for during the split into $\mathcal{D}_{\text{train}}$ and $\mathcal{D_{\text{new}}}$ or during the creation of CV splits on $\mathcal{D}_{\text{train}}$, resulting in a potential optimistic bias for both $\widehat{\mathrm{PE}}_{\text{new}}$ (due to the initial split) and $\widehat{\mathrm{PE}}_{\text{train}}$ (due to the CV splits). As $\widehat{\mathrm{PE}}_{\text{train}}$ is also subject to a larger clustering-induced optimistic bias than $\widehat{\mathrm{PE}}_{\text{new}}$, the bias does not cancel out when taking their difference and is therefore evident in Figure 5. Notably, the different levels of clustering-induced optimistic bias in $\widehat{\mathrm{PE}}_{\text{train}}$ and $\widehat{\mathrm{PE}}_{\text{new}}$ appear to have less impact on $R^{2}$, where, as described above, the prediction error differences are mostly negative. Further details on the impact of the clustering structure on the results, including an explanation of why it was not considered when performing the splits, are provided in Supplementary Section B.5.
@@
The additional source of optimistic bias introduced by the clustering structure of the data is also relevant when interpreting the prediction error differences in the analysis settings with HP tuning. While our primary focus here is on overlap-induced data leakage that arises since the observations used for the CV-based error estimation have already been seen during HP tuning (thus hindering the detection of potential overtuning), we have to consider that any observed optimistic bias may as well stem from clustering-induced data leakage. Consequently, we compare the prediction error differences in analysis settings with HP tuning to those in settings without tuning (where only clustering-induced data leakage is present) rather than directly comparing them to zero. Based on this assessment, the impact of overlap-induced data leakage on $\widehat{\mathrm{PE}}_{\text{train}}$ appears to be limited. This is particularly true for RMSE, where the CV error differences are generally comparable to those resulting from the I-no tuning procedure. For $R^{2}$, the median differences tend to be closer to zero compared to the I-no tuning procedure. In some analysis settings involving the smaller sample size and the CART algorithm, there is even a positive median difference (with the largest median difference of 0.018 observed in the setting where II-automated-PA is used in combination with the smaller sample size and the CART algorithm). Consequently, there appears to be a small overtuning effect that is not detected by the CV error due to overlap-induced data leakage. However, the median differences are too close to zero, and the variation within each analysis setting is too large to definitively determine which bias ultimately predominates, i.e. whether the CV error is overall optimistic or pessimistic in these settings.
@@
To summarize, the choice of model generation and evaluation procedure generally affects the difference between the prediction error estimates derived from available data and new data. As expected, when the evaluation procedure is based on the apparent error, the resulting estimate exhibits an optimistic bias, which varies depending on the model generation procedure. As likewise expected, the simple CV error is less optimistic than the apparent error, while the nested CV error is even less optimistic. The corresponding prediction error differences are less variable across model generation procedures compared to the apparent error. For simple CV, this indicates that, in the considered experimental setup, the tuning procedures do not introduce relevant overtuning effects on error estimation. Instead, the main source of bias for simple CV is either the clustering-induced optimistic bias (or, more precisely, the different bias level relative to $\widehat{\mathrm{PE}}_{\text{new}}$) or the pessimistic bias arising from the use of fewer observations during evaluation. This also holds true for the nested CV error.
@@ 6 Discussion and conclusion
To provide different examples of model generation and evaluation procedures in the context of HP tuning and to examine their impact on the difference between prediction error estimates from available and new data, we conducted an illustrative study using a real-world prediction problem from palliative care medicine. Although both the apparent error and CV error can, in theory, be optimistically biased when HPs are tuned, this was consistently true only for the apparent error (with the highest optimistic bias occurring in analysis settings that imitated manual tuning of preprocessing HPs without considering algorithm HPs). In contrast, the prediction error differences for the CV error appeared not to be considerably compromised by data leakage, as these differences were comparable to the analysis settings without HP tuning.
@@
Based on these conceptual and empirical insights, it is clear that to ensure HP tuning becomes a benefit rather than a pitfall, applied ML users must take care throughout the entire model development process. First, they should thoroughly consider which HPs (including preprocessing HPs) are to be tuned and which are not. An adequate tuning procedure that fits the specific prediction problem should then be specified. Unfortunately, this is typically non-trivial, as it depends on various factors such as sample size and the specific HPs to be tuned. More research is needed to better guide users in this respect (see [78], for an overview of current recommendations). In general, it is recommended to use automated tuning procedures instead of manual ones (see again [78], for automated tuning implementations in R and Python). If automated tuning is not feasible, users should at least ensure that the manual tuning procedure is error-free, reproducible, and resampling-based. For model evaluation, only two evaluation procedures are guaranteed to be unaffected by data leakage caused by HP tuning: (i) nested resampling (if the entire data set is used for model generation) or (ii) a permanent (outer) holdout (if only a subset of the available data is used for model generation). However, similar to the tuning procedure, there is a lack of guidance on how to choose between these approaches and how to specify them (e.g., which resampling methods to use for nested resampling). Although simple resampling may turn out to be a viable option in some applications (including our example), this can generally not be known in advance. Therefore, we discourage its use in settings involving HP tuning, as well as any other evaluation procedures that could result in data leakage.
Regardless of how model generation and evaluation are performed, it is essential that they and all other relevant details (e.g., the complete learning pipeline and its HPs) are transparently reported in both code and text form. For this purpose, users may rely on checklists such as REFORMS ([107]; intended for all applied research fields using ML) or TRIPOD+AI ([85]; intended for clinical prediction models). While transparency does not imply correctness, it allows readers to identify potential issues, such as data leakage, and to critically interpret the claimed model performance. Moreover, it emphasizes the existence and importance of preprocessing and its HPs, while the current lack of transparency can create the impression that the data were not preprocessed at all or that no alternative preprocessing options were explored. To further enhance transparency and encourage applied ML users to be more intentional about their choices, it is also possible to preregister the entire model development process, for example, by using the template proposed by [100].
@@
This work was supported by the German Research Foundation (BO3139/9-1, BO3139/7) to ALB. The authors of this work take full responsibility for its content.
@@ Appendix A Other leakage types
As stated in Section 2.4.2, [108] identify three general types of data leakage, which may arise from: (i) overlap between the data used for model generation and evaluation, (ii) violation of the assumption that all observations are independently drawn from the same distribution, or (iii) use of illegitimate features. While our paper primarily addresses overlap-induced data leakage, we will now provide additional details on the other two types.
@@
Even with a strict separation between the data used for model generation and evaluation, achieved through the use of resampling methods, data leakage can still occur if the assumption that all observations in $\mathcal{D}_{\text{train}}$ are independently drawn from the same distribution is violated. This assumption, also known as the i.i.d. assumption, was stated in Section 2.1. Non-i.i.d. settings may, for example, arise when $\mathcal{D}_{\text{train}}$ is a clustered data set, i.e. when the observations originate from different clusters (e.g., study centers). Observations within clusters are typically more similar than observations between clusters, where similarity can refer to both the feature vector $\boldsymbol{x}^{(i)}$ or the outcome $y^{(i)}$ [103]. If the prediction model is intended to be applied to observations from other clusters than those present in $\mathcal{D}_{\text{train}}$ in the future, resampling methods that are based on random sampling (i.e. ignoring the cluster structure) will be optimistically biased since in each resampling iteration, the observations in $\mathcal{D}_{\text{test}}$ are more similar to $\mathcal{D}^{\prime}_{\text{train}}$ than observations originating from new clusters [103, 108, 127]. Although the level of optimistic bias depends on the specific clustering structure (e.g., cluster size and correlation within clusters), it is generally recommended to perform grouped resampling at cluster level, where all observations in a cluster are either assigned to $\mathcal{D}^{\prime}_{\text{train}}$ or $\mathcal{D}_{\text{test}}$ in each resampling iteration [103, 78]. In the context of healthcare research, this type of resampling is referred to as internal-external validation [84, 87]. For other examples of non-i.i.d. settings and corresponding resampling methods, see [103] and the references therein.
@@
In Setting II, where resampling is typically used for both model generation (tuning) and evaluation, data leakage due to the violation of the i.i.d. assumption biases the prediction error estimate of the final model only when the non-i.i.d. data structure is ignored during model evaluation. This occurs specifically in the outer resampling loop of nested resampling (for $\mathcal{D}_{\text{train}}=\mathcal{D}$) or in the permanent outer holdout (for $\mathcal{D}_{\text{train}}\subset\mathcal{D}$). However, it is recommended to also take into account the non-i.i.d. data structure during tuning, both for the final prediction model and, if nested resampling is used, within the inner resampling loop, to ensure consistency [103].
@@ A.2 Use of illegitimate features
If $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ include features that are generally not available for new observations to which the model will be applied in practice, these features can be considered illegitimate, and if included in the final prediction model, constitute another type of data leakage. An example raised by [108] is the use of anti-hypertensive drugs as a feature for predicting hypertension. Note that this type of data leakage is conceptually different from the other two types, as it stems from a design issue that is independent of the model evaluation procedure.
@@ B.1 Descriptive statistics
Table S1: Distribution of the outcome variable and features in the COMPANION data set after applying the initial preprocessing steps (described in Supplementary Section B.2.1). In addition, two preprocessing steps from the learning pipeline $\mathcal{I}$ (see Section 5.2.2 and Supplementary Section B.2.2) have been performed: the correction of costs and the aggregation of the IPOS score (default version).
n=n=1,449